OpenStudy (anonymous):

The lengths of two sides of a right triangle are given. Find the length of the third side. Round to the nearest tenth if necessary. legs: 28 in. and 15 in. A. 37.5 in. B. 23.6 in. C. 29.6 in. D. 31.8 in.

OpenStudy (bahrom7893):

pythagorean theorem

OpenStudy (bahrom7893):

=sqrt(28^2+15^2)

OpenStudy (hoblos):

x^2 = 28^2 + 15^2 = 1009 x= sqrt(1009) = 31.8 in

OpenStudy (anonymous):

D. 31.8 in sqrt28 + sqrt15 = 1009 the square root of 1009 is approx. 31.8